if the sum of four numbers in A.P is24 and sum of thier squares is 164.then those numbers are

Grade:11

2 Answers

Kaushik
10 years ago

they are 3,5,7,9

srinivas aytha
10 years ago

let the four no be a-3d,a-d,a+d,a+3d.

a-3d+a-d+a+d+a+3d=24 (given)

4a=24

a=6.

(a-3d)2+(a-d)2+(a+d)2+(a+3d)2=164 (given)

2(a2+9d2)+2(a2+d2)=164. since (a+b)2+(a-b)2=2(a2+b2)

2(2a2+10d2)=164.

2a2+10d2=164/2=82.

2(6)2+10d2=82.

10d2=10.

d=±1.

therefore four numbers are 3,5,7,9